Bella and Regina both leave the mall at the same time, but in opposite directions. If Regina travels 5 mph faster than Bella and after 3 hours they are 87 miles apart, how fast is each traveling?

Found 3 solutions by ikleyn, math_tutor2020, Alan3354:
Answer by ikleyn(52777)

.
Bella and Regina both leave the mall at the same time, but in opposite directions.
If Regina travels 5 mph faster than Bella and after 3 hours they are 87 miles apart,
how fast is each traveling?
~~~~~~~~~~~~~~~~~


Let r be the slower rate;  then the faster rate is (r+5) mph.


The total separating distance in 3 hours is  3r + 3*(r+5) miles,

and it is equal to 87 miles.  So your base equation is

    3r + 3*(r+5) = 87  miles.


Simplify and find r

    3r + 3r + 15 = 87

         6r      = 87 - 15 = 72

         r                 = 72/6 = 12 miles per hour.


ANSWER.  The Bella's rate is   12 mph.

         The Regina rate is  12 + 5 = 17 mph.


CHECK.  3*12 + 3*17 = 36 + 51 =  87  miles separating distance in 3 hours.   ! Correct !

Solved.

-----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.



Answer by math_tutor2020(3816)


B = Bella's speed in mph
B+5 = Regina's speed in mph

The formula we'll need is:
distance = rate*time
often abbreviated as
d = r*t
The term "rate" is the same as "speed".

If Bella travels for 3 hours at a speed of B mph, then she travels a distance of...
d = r*t
d = B*3
d = 3B
If we knew what B was, then we could find a numeric distance for Bella.

Meanwhile, Regina also travels for 3 hours but at a speed of B+5 mph
d = r*t
d = (B+5)*3
d = 3B+15

Here are their distances:
Bella = 3B
Regina = 3B+15
These distances are in miles.

These two distances must add to 87 miles since this is the gap between them after the three hour mark.
It might help to draw out a number line diagram to visualize what's going on.
Have one person go left, and the other go right.

Bella + Regina = 87
(3B) + (3B+15) = 87
6B + 15 = 87
6B = 87-15
6B = 72
B = 72/6
B = 12
Bella's speed is 12 mph.
This is comparable to bike speed, though a bit on the slower side.
Perhaps Bella is riding up a slightly steep hill, or she may be biking at a somewhat leisurely pace not in a hurry to get anywhere.

Regina's speed is B+5 = 12+5 = 17 mph

If you wanted, you can calculate the distances of each person (now that we've determined the value of B). I'll leave that to you for practice.

Answers:
Bella's speed = 12 mph
Regina's speed = 17 mph

Answer by Alan3354(69443)

"...both leave the mall at the same time..."
------------------
As opposed to only one of them leaving at the same time?